OpenStudy (anonymous):

Find x in [−6π, 0] such that cos(x/6) = 1/2

OpenStudy (mimi_x3):

\[cos\left(\frac{x}{6}\right) =\frac{1}{2}\] well have you tried it yet

OpenStudy (anonymous):

yes. it doesnt work

OpenStudy (mimi_x3):

what have you tried

OpenStudy (anonymous):

i tried doing cos^-1(1/2) and multiplying by 6

OpenStudy (mimi_x3):

do you know general solutions?

OpenStudy (anonymous):

no haha

OpenStudy (mimi_x3):

\[\cos\left(\frac{x}{6}\right) =\frac{1}{2} => \frac{x}{6} =\cos^{-1}\left(\frac{1}{2}\right) => \frac{x}{6} =\frac{\pi}{3} \]

OpenStudy (mimi_x3):

are you able to do it now

OpenStudy (anonymous):

no that didnt work :( but i feel like thats almost right. How did you know it equaled pi/3?

OpenStudy (mimi_x3):

\[\frac{x}{6} =\frac{\pi}{3} => 3x = 6\pi => x= 2\pi \] it works

OpenStudy (anonymous):

i didnt accept it lol